OpenStudy (explainitlikeimfive):

The area of a rectangle is x2 - 2x - 15 and the length of the rectangle is x + 3. Find the width of the rectangle. A. x – 5 B. x2 + 15 C. x + 5 D. (x – 2)(x + 3)

jimthompson5910 (jim_thompson5910):

Hint: divide the area by the length and factor x^2 - 2x - 15 as much as possible

OpenStudy (explainitlikeimfive):

x2 - 2x - 15 = (x+3)(x-5)

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

now divide that by the length

jimthompson5910 (jim_thompson5910):

you'll see cancellations

jimthompson5910 (jim_thompson5910):

this works because Area = Length*Width A = L*W A/L = W W = A/L

OpenStudy (explainitlikeimfive):

I end up with \[\frac{ 5 }{ x }\]

jimthompson5910 (jim_thompson5910):

that's not correct

OpenStudy (explainitlikeimfive):

I re-did it and still got the same answer. Im doing something wrong

jimthompson5910 (jim_thompson5910):

(x+3)(x-5) divided by x+3 is just x-5

jimthompson5910 (jim_thompson5910):

since the x+3 terms cancel

OpenStudy (explainitlikeimfive):

That makes sense. Thank you very much.

jimthompson5910 (jim_thompson5910):

yw